# “Hello, is this Anna?": A First Look at Pig-Butchering Scams

### v1

@@ 4.6 The Cut
Stalling Tactics. The cut phase typically begins when victims refuse to invest any more money, and attempt to withdraw funds. Scammers used several tactics to delay the withdrawal; 21 participants (80.8%) discovered that they could suddenly no longer access their investment accounts. 10 participants (38.4%) reported that the platforms displayed error messages or indicated that the account was under review. Some scammers went a step further; 16 participants reported that the platform asked for more money under the guise of withdrawal fees or taxes. 4 participants (15.4%) fell for this and ended up paying. P19 says:
@@
Complete Disappearance. After either paying fees for withdrawal, or refusing to do so, all but one participants reported the scammer ceasing all communication. Scammers blocked the victims on chat application (WhatsApp / Telegram), as well as any social media platforms they were connected on. Scammer accounts were likely deleted or deactivated; 18 participants (69.2%) searched for the profiles using other accounts, but could not find them. P22 says:
@@ 4.7 The Encore
Impersonation Scams. Scammers frequently impersonate law enforcement officials to re-establish contact with victims. They pose as detectives, federal agents, or members of cybercrime units investigating the original scam. 15 participants (57.7%) reported receiving official-looking communications, complete with badges, case numbers, and legal jargon. 8 (30.7%) reported that the supposed agent requested personal information under the guise of needing it for the investigation. According to P12:
@@
Further, 9 participants (34.6%) reported that these fraudulent officials informed them that their funds have been located but require payment of legal fees or taxes to release them. Overall, 15 participants were victims of impersonation scams, and ended up paying additional amounts of \$120 – \$1700.
Recovery Scams. Scammers, sometimes the original perpetrators or new opportunists, contact victims offering services to retrieve lost funds. 11 participants reported being contacted by such recovery experts. They present themselves as recovery specialists ($n=7$, 26.9%) private investigators ($n=2$, 7.6%), or legal professionals with expertise in financial fraud cases ($n=5$, 19.2%). They typically request upfront fees for their services, citing costs for legal processes, court fees, or administrative expenses. Overall, 11 participants were approached with recovery scam offers. These scams can be highly sophisticated; P20 says:
@@
Legal Threats. Scammers may also employ legal threats to intimidate victims into making additional payments. They impersonate lawyers ($n=4$, 15.4%), debt collectors ($n=1$, 3.8%), or officials from regulatory agencies ($n=2$, 8.6%), claiming that the victim has breached contracts or violated laws through their previous investment activities. The threats typically include demands for immediate payment to avoid lawsuits, arrests, or damage to credit ratings. Scammers may threaten to report the victim to authorities for participating in illegal investment schemes or non-payment of taxes. P21 reported:
@@
Overall, 7 participants reported being contacted with legal threats; however, fortunately, none of them fell for it. {mdframed}[style=insightstyle] In some cases, scammers re-engage victims by posing as law enforcement or recovery agents, offering assistance in recovering lost funds for a fee. This phase exploits the victim’s emotional vulnerability and desperation, further compounding their financial losses.
Phase Duration Strategies Used
Approached via dating apps ($n=6$, 23%)
Approached via social media ($n=2$, 7.7%)
Developed business partnerships ($n=10$, 38.5%)
Shared personal stories ($n=8$, 30.7%)
Discussed family and career ($n=11$, 42.3%)
Showed empathy ($n=5$, 19.2%)
Mirroring manipulation ($n=9$, 34.6%)
Sent frequent pictures ($n=17$, 65.4%)
Audio calls ($n=7$, 26.9%)
Video calls ($n=3$, 11.5%)
Introduced crypto as hobby ($n=3$, 11.5%)
Claimed insider knowledge ($n=19$, 73.1%)
Investment consultant managing portfolio ($n=7$, 26.9%)
Framed as exclusive opportunity ($n=6$, 23%)
Provided fake profit screenshots ($n=22$, 84.6%)
Showed messages confirming returns ($n=5$, 19.2%)
Provided contacts of supposed investors ($n=4$, 15.4%)
Emphasized exclusivity and FOMO ($n=18$, 69.2%)
Assisted in account setup ($n=11$, 42.3%)
Platforms had help/support ($n=5$, 19.2%)
Allowed small withdrawals ($n=3$, 11.5%)
Encouraged reinvestment of gains ($n=23$, 88.5%)
Suggested larger investments ($n=13$, 50%)
Manipulated platform to appear co-investing ($n=14$, 53.8%)
Introduced complex strategies ($n=8$, 30.7%)
Discouraged discussing with others
Claimed insider deals access ($n=12$, 46.2%)
Claimed ICOs about to explode ($n=9$, 34.6%)
Increased urgency of communication (all participants)
Emotional manipulation and disappointment ($n=11$, 42.3%)
Encouraged using credit to invest ($n=9$, 34.6%)
Accounts under review/errors ($n=10$, 38.5%)
Requested fees/taxes for withdrawal ($n=16$, 61.5%)
Ceased all contact ($n=25$, 96.1%)
Scammer profiles disappeared ($n=18$, 69.2%)
Requested personal info for investigation ($n=8$, 30.8%)
Claimed funds found, need fees/taxes ($n=9$, 34.6%)
Recovery scams by specialists ($n=7$, 26.9%)
Recovery scams by private investigators ($n=2$, 7.7%)
Recovery scams by legal professionals ($n=5$, 19.2%)
Legal threats by impersonated lawyers ($n=4$, 15.4%)
Legal threats by debt collectors ($n=1$, 3.8%)
Legal threats by regulatory officials ($n=2$, 7.7%)
Table 1: Summary of Pig-Butchering Scam Lifecycle
@@ 5.1 Financial Impact
Loss of Savings. A significant number of victims experienced devastating losses to their personal savings. Many had invested substantial portions of their life savings, retirement funds, or college funds for their children, enticed by the promise of high returns. Financial strain affected victims’ ability to meet daily living expenses, and resulted in long-term financial insecurity. 15 participants (57.7%) reported struggling to pay for rent, medical care or other bills. 6 participants (who were retired or nearing retirement age) had used their retirement savings to invest; P22 says:
@@
Accumulation of Debt. Many victims accrued substantial debt as a result of the scam. 12 participants (46.2%) resorted to personal loans, borrowing from friends and family, and 4 (15.4%) liquidated an existing asset to invest more heavily under the scammer’s influence. In the words of P25:
@@
Further, 13 participants (50%) reported accruing significant credit card debt. High-interest rates associated with credit cards further exacerbated their financial difficulties post-scam. 2 (7.7%) even reported damaged credit scores due to missed payments and increased debt-to-income ratios.
Ripple Effect. Finally, for 5 participants (19.23%), financial losses affected their dependents and family members significantly. P16 recounts:
@@ 5.2 Emotional Impact
Shame and Embarassment. The most common feelings experienced by victims were shame and embarrassment, as reported by 19 participants (73%). P3 expressed:
@@
Feelings of stupidity and self-blame were also common, as victims grappled with how they could have been manipulated. 10 participants (38.5%) reported that they blamed themselves for falling for something that was obviously a scam. The stigma associated with being scammed heightened their sense of humiliation; P25 mentioned:
@@
Loss of Trust. Loss of trust was another critical emotional impact; betrayal by scammers severely impacted victims’ ability to trust others. According to 22 participants (84.6%), it was hard to trust anyone online, even if they were previously known to them, fearing that a scammer was using a fake profile or hijacked account. Participants who developed romantic relationships during the scam reported experiencing feelings akin to heartbreak; P26 says:
@@
Emotional Distress & Mental Health. The emotional distress caused by the scam had significant implications for victims’ mental health. Many reported being clinically diagnosed with mental health conditions such as depression ($n=4$, 15.4%), anxiety ($n=9$, 34.6%), and even post-traumatic stress disorder (PTSD) ($n=2$, 7.7%). Paranoia and hyper-vigilance were also reported; 8 participants (30.8%) described themselves as being constantly on edge, fearing further exploitation. P6 mentioned:
@@ 5.3 Seeking Help from Law Enforcement
Reluctance to Report. A significant number of participants ($n=19$, 73%) did not report the scam to authorities due to feelings of shame and embarrassment. Feelings of shame and self-blame deterred many from coming forward. In the words of P7:
@@
The stigma associated with being a victim of fraud contributed to a culture of silence; according to 7 participants (26.9%), they did not want to come forward with their stories because they did not want anyone to know that they had been a victim. P8 said:
@@
Limited Assistance from Authorities. Those who did report the scam often found that law enforcement could do little to assist. 7 participants reported lodging a complaint with their local police station, and 3 with the FBI. Out of the 7 participants who sought help from their local police, 6 were informed that, without formal contracts or identifiable perpetrators, pursuing the case was challenging. 2 were also told that the transactions weren’t fraudulent; P21 mentioned:
@@ 5.4 Pig Butchering Terminology
The term "pig butchering" elicited strong emotional reactions from the victims interviewed in this study. Most participants ($n=23$, 88.5%) expressed discomfort and offense at being associated with the term, feeling that it dehumanized their experiences and added insult to injury.P7 remarked:
@@
18 victims (69.2%) felt that the term trivialized their ordeal, reducing a complex and traumatic experience to a crude metaphor. In the words of P11:
@@
Further, 17 participants (65.4%) also felt that the term was mocking, and that that was one of the contributing factors in them not reporting the scam. P13 stated:
@@ 6.2 Implications
Law Enforcement. For law enforcement, the complexity and anonymity of pig butchering scams pose a significant challenge, as victims reported high levels of frustration with the response they received. Strengthening cybercrime units with specialized training and resources dedicated to tackling financial scams is essential to improve response effectiveness. Law enforcement agencies could benefit from establishing victim-focused units trained in cyber psychology and financial fraud to ensure sensitive handling of such cases. Improved digital forensic capabilities would enable agencies to track scammers more effectively, even across borders. Additionally, law enforcement agencies should coordinate with banks and financial institutions to monitor unusual transactions that could signal fraud. Cross-agency collaboration between cybercrime units and financial regulators could help streamline response times and increase scam intervention success rates. Given the stigma victims feel, law enforcement should also create accessible, stigma-free channels for victims to report scams anonymously.
Government Policy & International Cooperation. The cross-border nature of pig butchering scams, with perpetrators frequently operating from regions with limited extradition agreements, underscores the need for governments to prioritize international cooperation in cybercrime investigations. These scams are often highly organized, requiring governments to strengthen cross-border frameworks that facilitate information sharing, investigation, and prosecution. Current anti-fraud laws are not uniformly enforceable internationally, which limits the ability to pursue justice for victims. Governments could support dedicated cybercrime units and establish clear frameworks for joint operations with other countries to improve response rates. Awareness campaigns run by governments can also play a vital role in educating the public on these scams, emphasizing their sophisticated emotional and psychological aspects. Through a concerted policy focus, governments can better protect citizens from becoming victims of international fraud schemes.
Rebranding Pig-Butchering. Replacing the term "pig-butchering" with more neutral, non-stigmatizing terminology could encourage victims to come forward and report their experiences. Adopting terminology that accurately describes the manipulative tactics without demeaning victims can reduce the stigma associated with reporting and discussing the scam. For example, terms like Grooming Investment Scam, Long-Con Investment Fraud or Groom-and-Swindle Scheme capture the essence of the scam in a non-derogatory way.
@@ Conclusion
Appendix D Participant Demographics
The demographics of our participants are depicted in Table LABEL:tab:demo.

### v2

@@ 4.6 The Cut
Stalling Tactics. The cut phase typically begins when victims refuse to invest any more money, and attempt to withdraw funds. Scammers used several tactics to delay the withdrawal; most participants discovered that they could suddenly no longer access their investment accounts. 10 participants reported that the platforms displayed error messages or indicated that the account was under review. Some scammers went a step further; 16 participants reported that the platform asked for more money under the guise of withdrawal fees or taxes. 4 participants fell for this and ended up paying. P19 says:
@@
Complete Disappearance. After either paying fees for withdrawal, or refusing to do so, all but one participants reported the scammer ceasing all communication. Scammers blocked the victims on chat application (WhatsApp / Telegram), as well as any social media platforms they were connected on. Scammer accounts were likely deleted or deactivated; participants searched for the profiles using other accounts, but could not find them. P22 says:
@@ 4.7 The Encore
Impersonation Scams. Scammers frequently impersonate law enforcement officials to re-establish contact with victims. They pose as detectives, federal agents, or members of cybercrime units investigating the original scam. Most participants reported receiving official-looking communications, complete with badges, case numbers, and legal jargon. Some reported that the supposed agent requested personal information under the guise of needing it for the investigation. According to P12:
@@
Further, a few participants reported that these fraudulent officials informed them that their funds have been located but require payment of legal fees or taxes to release them. Overall, 15 participants were victims of impersonation scams, and ended up paying additional amounts of \$120 – \$1700.
Recovery Scams. Scammers, sometimes the original perpetrators or new opportunists, contact victims offering services to retrieve lost funds. 11 participants reported being contacted by such recovery experts. They present themselves as recovery specialists, private investigators, or legal professionals with expertise in financial fraud cases . They typically request upfront fees for their services, citing costs for legal processes, court fees, or administrative expenses. Overall, 11 participants were approached with recovery scam offers. These scams can be highly sophisticated; P20 says:
@@
Legal Threats. Scammers may also employ legal threats to intimidate victims into making additional payments. They impersonate lawyers, debt collectors, or officials from regulatory agencies, claiming that the victim has breached contracts or violated laws through their previous investment activities. The threats typically include demands for immediate payment to avoid lawsuits, arrests, or damage to credit ratings. Scammers may threaten to report the victim to authorities for participating in illegal investment schemes or non-payment of taxes. P21 reported:
@@
{mdframed}
[style=insightstyle] In some cases, scammers re-engage victims by posing as law enforcement or recovery agents, offering assistance in recovering lost funds for a fee. This phase exploits the victim’s emotional vulnerability and desperation, further compounding their financial losses.
@@ 5.1 Financial Impact
Loss of Savings. A significant number of victims experienced devastating losses to their personal savings. Many had invested substantial portions of their life savings, retirement funds, or college funds for their children, enticed by the promise of high returns. Financial strain affected victims’ ability to meet daily living expenses, and resulted in long-term financial insecurity. Some reported struggling to pay for rent, medical care or other bills. 6 participants (who were retired or nearing retirement age) had used their retirement savings to invest; P22 says:
@@
Accumulation of Debt. Many victims accrued substantial debt as a result of the scam. A few resorted to personal loans, borrowing from friends and family, and liquidated an existing asset to invest more heavily under the scammer’s influence. In the words of P25:
@@
Further, participants reported accruing significant credit card debt. High-interest rates associated with credit cards further exacerbated their financial difficulties post-scam. Some even reported damaged credit scores due to missed payments and increased debt-to-income ratios.
Ripple Effect. Finally, for 5 participants, financial losses affected their dependents and family members significantly. P16 recounts:
@@ 5.2 Emotional Impact
Shame and Embarassment. The most common feelings experienced by victims were shame and embarrassment, as reported by most participants. P3 expressed:
@@
Feelings of stupidity and self-blame were also common, as victims grappled with how they could have been manipulated. A few reported that they blamed themselves for falling for something that was obviously a scam. The stigma associated with being scammed heightened their sense of humiliation; P25 mentioned:
@@
Loss of Trust. Loss of trust was another critical emotional impact; betrayal by scammers severely impacted victims’ ability to trust others. According to most participants, it was hard to trust anyone online, even if they were previously known to them, fearing that a scammer was using a fake profile or hijacked account. Participants who developed romantic relationships during the scam reported experiencing feelings akin to heartbreak; P26 says:
@@
Emotional Distress & Mental Health. The emotional distress caused by the scam had significant implications for victims’ mental health. Many reported being clinically diagnosed with mental health conditions such as depression, anxiety, and even post-traumatic stress disorder (PTSD). Paranoia and hyper-vigilance were also reported; some described themselves as being constantly on edge, fearing further exploitation. P6 mentioned:
@@ 5.3 Seeking Help from Law Enforcement
Reluctance to Report. A significant number of participants did not report the scam to authorities due to feelings of shame and embarrassment. Feelings of shame and self-blame deterred many from coming forward. In the words of P7:
@@
The stigma associated with being a victim of fraud contributed to a culture of silence; according to 7 participants, they did not want to come forward with their stories because they did not want anyone to know that they had been a victim. P8 said:
@@
Limited Assistance from Authorities. Those who did report the scam often found that law enforcement could do little to assist. Only 7 participants reported lodging a complaint with their local police station, and 3 with the FBI. Out of those who sought help from their local police, 6 were informed that, without formal contracts or identifiable perpetrators, pursuing the case was challenging. 2 were also told that the transactions weren’t fraudulent; P21 mentioned:
@@ 5.4 Pig Butchering Terminology
The term "pig butchering" elicited strong emotional reactions from the victims interviewed in this study. Most participants expressed discomfort and offense at being associated with the term, feeling that it dehumanized their experiences and added insult to injury.P7 remarked:
@@
Most also felt that the term trivialized their ordeal, reducing a complex and traumatic experience to a crude metaphor. In the words of P11:
@@
Further, participants also felt that the term was mocking, and that that was one of the contributing factors in them not reporting the scam. P13 stated:
@@ 6.2 Implications
Law Enforcement. For law enforcement, the complexity and anonymity of pig butchering scams pose a significant challenge, as victims reported high levels of frustration with the response they received. Strengthening cybercrime units with specialized training and resources dedicated to tackling financial scams is essential to improve response effectiveness. Law enforcement agencies could benefit from establishing victim-focused units trained in cyber psychology and financial fraud to ensure sensitive handling of such cases. Improved digital forensic capabilities would enable agencies to track scammers more effectively, even across borders. Additionally, law enforcement agencies should coordinate with banks and financial institutions to monitor unusual transactions that could signal fraud. Cross-agency collaboration between cybercrime units and financial regulators could help streamline response times and increase scam intervention success rates.
Government Policy & International Cooperation. The cross-border nature of pig butchering scams, with perpetrators frequently operating from regions with limited extradition agreements, underscores the need for governments to prioritize international cooperation in cybercrime investigations. These scams are often highly organized, requiring governments to strengthen cross-border frameworks that facilitate information sharing, investigation, and prosecution. Current anti-fraud laws are not uniformly enforceable internationally, which limits the ability to pursue justice for victims. Governments could support dedicated cybercrime units and establish clear frameworks for joint operations with other countries to improve response rates. Awareness campaigns run by governments can also play a vital role in educating the public on these scams, emphasizing their sophisticated emotional and psychological aspects.
Rebranding Pig-Butchering. Replacing the term "pig-butchering" with more neutral, non-stigmatizing terminology could encourage victims to come forward and report their experiences. Adopting terminology that accurately describes the manipulative tactics without demeaning victims can reduce the stigma associated with reporting and discussing the scam. For example, terms like Grooming Investment Scam, Long-Con Investment Fraud or Groom-and-Swindle Scheme capture the essence of the scam in a non-derogatory way. We still use the term “pig-butchering" because at the time of writing, it was the most widely recognized term in public discourse, law enforcement advisories, and emerging literature. By first anchoring the discussion using the recognizable term, and then problematizing it with evidence from victim narratives, we believe we can effectively advocate for re-framing in the terminology.
@@ Conclusion
Appendix D Detailed Lifecycle of Pig-Butchering Scams
To synthesize our findings, we present Table 1, which outlines the seven-stage lifecycle of pig-butchering scams as reported by victims in our study. Each stage captures a distinct phase in the scam’s progression—from initial contact to emotional manipulation, financial exploitation, and post-scam re-engagement.
Phase Duration Strategies Used
Approached via dating apps ($n=6$, 23%)
Approached via social media ($n=2$, 7.7%)
Developed business partnerships ($n=10$, 38.5%)
Shared personal stories ($n=8$, 30.7%)
Discussed family and career ($n=11$, 42.3%)
Showed empathy ($n=5$, 19.2%)
Mirroring manipulation ($n=9$, 34.6%)
Sent frequent pictures ($n=17$, 65.4%)
Audio calls ($n=7$, 26.9%)
Video calls ($n=3$, 11.5%)
Introduced crypto as hobby ($n=3$, 11.5%)
Claimed insider knowledge ($n=19$, 73.1%)
Investment consultant managing portfolio ($n=7$, 26.9%)
Framed as exclusive opportunity ($n=6$, 23%)
Provided fake profit screenshots ($n=22$, 84.6%)
Showed messages confirming returns ($n=5$, 19.2%)
Provided contacts of supposed investors ($n=4$, 15.4%)
Emphasized exclusivity and FOMO ($n=18$, 69.2%)
Assisted in account setup ($n=11$, 42.3%)
Platforms had help/support ($n=5$, 19.2%)
Allowed small withdrawals ($n=3$, 11.5%)
Encouraged reinvestment of gains ($n=23$, 88.5%)
Suggested larger investments ($n=13$, 50%)
Manipulated platform to appear co-investing ($n=14$, 53.8%)
Introduced complex strategies ($n=8$, 30.7%)
Discouraged discussing with others
Claimed insider deals access ($n=12$, 46.2%)
Claimed ICOs about to explode ($n=9$, 34.6%)
Increased urgency of communication (all participants)
Emotional manipulation and disappointment ($n=11$, 42.3%)
Encouraged using credit to invest ($n=9$, 34.6%)
Accounts under review/errors ($n=10$, 38.5%)
Requested fees/taxes for withdrawal ($n=16$, 61.5%)
Ceased all contact ($n=25$, 96.1%)
Scammer profiles disappeared ($n=18$, 69.2%)
Requested personal info for investigation ($n=8$, 30.8%)
Claimed funds found, need fees/taxes ($n=9$, 34.6%)
Recovery scams by specialists ($n=7$, 26.9%)
Recovery scams by private investigators ($n=2$, 7.7%)
Recovery scams by legal professionals ($n=5$, 19.2%)
Legal threats by impersonated lawyers ($n=4$, 15.4%)
Legal threats by debt collectors ($n=1$, 3.8%)
Legal threats by regulatory officials ($n=2$, 7.7%)
Table 1: Summary of Pig-Butchering Scam Lifecycle